### v1

@@ 4.2 Model statistics and extreme value estimates of water level
The areas that are most influenced by storm surges in the area covered by our hindcast are the coast of the United Kingdom, the Netherlands, the West Coast of Denmark, the area around Lofoten in Norway and the Russian coast between the Kola Peninsula and Novaya Zemlya. Generally, as shown in Figure 8, these areas contain both the highest and lowest values of storm surge in the hindcast period, and hence also the largest ranges of storm surge. The largest minimum value for storm surge in the hindcast, as can be seen in Figure 8a, is $-2.60~{}m$, and is found on the south eastern coast of the UK, near the mouth of the river Thames. The largest value for maximum storm surge, as shown in Figure 8b, is $3.92~{}m$, and is located in the German Bight near the mouth of the river Elbe close to Hamburg in Germany. This area is a hot spot for storm surges, and also includes the largest storm surge range in the hindcast of $5.65~{}m$, as seen in Figure 8c. Generally, the southeastern part of the North Sea is an area of large storm surge activity, as is also illustrated by the large values of standard deviation for the DE, NL and DK stations in Figure 3a.
@@
We acknowledge the fact that a 43-year long hindcast is not a sufficiently long period of time to provide robust statistics of extreme water levels by direct analysis and extraction of extremes from the model results. We have therefore performed a statistical analysis based on annual (block) maxima of the hindcast dataset and calculated the 100-year Return Value Estimates (RVE) of minimum and maximum storm surge for the model domain by fitting a Gumbel distribution (Gumbel, 1958) to the annual minima and maxima, respectively. As presented by Coles (2001) the cumulative distribution function (CDF) of the block maxima formed from a random sequence of independent variables and identically distributed (IID) variables will follow the Generalized Extreme Value (GEV) distribution,
@@ 4.3.1 The 1987 Oslofjord storm surge
As the low pressure system moved further east, and the wind direction continued backing further left, the storm surge signal moved further through the North Sea up along the western coast of Denmark and into the Skagerrak area as a Kelvin wave (see Figure 12). Even though not directly affected by the weather system itself, the stations in the Oslofjord experienced large fluctuations in water level due to the Kelvin wave, with variations in water level due to storm surge in the inner part of the Oslofjord in the range of up to ±$\pm 1~{}m$ in less than 24 hours. This can be seen in the lower panel in Figure 12, displaying the observed and modeled storm surge for the station Viker (NO_VIK) at the entrance to the Oslofjord. Both the amplitude of the surge, and the values of MAE and RMSE, is significantly reduced compared to the two panels above for Cuxhaven and Esbjerg (DK_6401). Correlation between hindcast and observations for the three stations for the time period shown is almost identical at $0.93$ and $0.94$.
@@ 5 Summary and concluding remarks
The NORA-Surge hindcast is an open and freely available high-resolution storm surge hindcast based on a barotropic setup of the ROMS ocean model with 4 km horizontal resolution covering the areas of the North Sea, the Norwegian Sea and the Barents Sea. It includes the entire coast of the United Kingdom, the Netherlands and Norway (including Svalbard), in addition there is partial coverage of the coast of Denmark and western coast of Sweden. The temporal coverage of the hindcast is more than four decades, covering the time period from 1979 throughout 2022.
The hindcast has been extensively validated against an observational dataset of quality checked observations of water level from more than 90 coastal and offshore locations, as described in Section 2. Comparison of the hindcast with all the observational data yields average values of MAE and RMSE of $9.72$ and $12.39$ cm respectively (see Table 1 and 2), and correlation coefficients for each station primarily in the range of $0.80-0.93$ (see Figure 6). The absolute error varies between the stations, and has larger values for some geographical areas than others. The areas with the largest amplitudes in storm surge are also associated with the largest errors (see Figure 3 and Table 2). However, as shown in the Taylor diagram in Figure 6, the relative errors for the stations (with a few exceptions) are quite similar. There is a slight clustering of the stations from different countries, which reflects geographical and bathymetry specificities of the coastline depending on the location, but all the clusters have at least some overlap with each others.
Statistics of extreme values of minimum, maximum and range for storm surge within the hindcast are calculated and presented in Figure 8. Also, a return value estimation has been carried out based on the hindcast dataset, and the 100-year return value estimates of minima and maxima of water level for the entire hindcast model domain are presented in Figure 9.
@@
We acknowledge that the present hindcast, even when shown to possess realistic properties, could be improved in the future. As has been mentioned, increased horizontal resolution with more detailed bottom bathymetry and a more realistic coastline, would likely reduce some errors in the hindcast. In addition, the inclusion of tides should be considered to improve the quality, especially in regions with large tidal amplitudes and shallow water depths. Finally, we expect the atmospheric forcing in the storm surge model could be improved by including wave effects in the atmospheric momentum fluxes (Staneva et al., 2017; Bonaduce et al., 2020).
For users utilizing the hindcast dataset, we encourage to subtract the hindcast mean storm surge water level for each grid point from the data before use, as we have done. Additionally, various methods for correction could be explored and applied to the hindcast dataset to further improve the quality. But this was not within the scope of the present work.
@@ Acknowledgments
The license agreements of the water level observations we gathered do not allow to redistribute our dataset of in-situ observations. However, the dataset can be shared upon reasonable request and collaboration by taking contact with the corresponding author of this study. Moreover, the list of data sources presented in Table 3 can be used to retrieve again the data. Though this involves a significant amount of work, we provide a series of code repositories [will be made public upon acceptance of the manuscript for publication] that can be used as a basis or inspiration for building the aggregated dataset:
Appendix B: Observation data collection and quality control
Prefix
Source
Data request URL
NO
Kartverket - Norwegian Mapping Authority
https://api.sehavniva.no/tideapi_en.html
WL
Norwegian Meteorological Institute, oil platforms sensors
MetNo internal data: address data request to ocean@met.no
SW
Swedish Meteorological and Hydrological Institute
https://www.smhi.se/data/oceanografi/ladda-ner-oceanografiska-observationer/#param=sealevelrh2000,stations=all
DK
Danish Coastal Authority
https://kyst.dk/hav-og-anlaeg/maalinger-og-data/vandstandsmaalinger
DE
BFG - Federal Institute of Hydrology
https://www.bafg.de/EN/06_Info_Service/01_WaterLevels/waterlevels_node.html
NL
Rijkswaterstaat environment
https://waterinfo.rws.nl/#/publiek/waterhoogte
UK
BODC - British Oceanographic Data Center
https://www.bodc.ac.uk/data/hosted_data_systems/sea_level/uk_tide_gauge_network/processed/
@@
  • For each data source, a custom parser (or even several parsers, as some of the data sources have format differences between different time periods and stations) was developed to convert the different data formats used into a list of tuples:
    $(time\_of\_observation,water\_level\_observation,quality\_flag).$
    From there, a time series of the raw collected observation data is established for each station (note that the quality_flag is available on a data source dependent basis).
  • The time series for each station is, at this stage, a series of observation tuples, which temporal resolution varies between stations (and for some stations, with time), from 30 minutes (coarsest observations assimilated in our dataset) to 10 minutes (finest observations available for some stations). We perform a simple quality checking, and we first reject data that have either a faulty quality flag (when provided), or that have clearly faulty values (unrealistic absolute value). We then perform interpolation of these data onto a common time basis, with 10 minutes resolution (corresponding to UTC minutes equal to 0 modulo 10).
At this stage, consistent time series on a common time basis are obtained for all the stations. We then perform a second quality check, that is based on human investigation. In particular, the time series for each station was visually inspected for obvious problems such as large average trends in the data (this happens, for example, for the oil platforms, that are standing on the seafloor and subject to seabed subsidence), and change of the reference levels following sensor replacement, maintenance, or re-positioning. These errors are removed, and the time series are made into 0-mean data and manually corrected, either by linearly detrending the data, or computing the mean water elevation to subtract on segments of data in between sensor servicing.

### v2

@@ 4.2 Model statistics and extreme value estimates of water level
The areas that are most influenced by storm surges in the area covered by our hindcast are the coast of the United Kingdom, the Netherlands, the West Coast of Denmark, the area around Lofoten in Norway and the Russian coast between the Kola Peninsula and Novaya Zemlya. Generally, as shown in Figure 8, these areas contain both the highest and lowest values of storm surge in the hindcast period, and hence also the largest ranges of storm surge. The largest minimum value for storm surge in the hindcast, as can be seen in Figure 8a, is $-2.60~{}m$, and is found on the south eastern coast of the UK, near the mouth of the river Thames. The largest value for maximum storm surge, as shown in Figure 8b, is $3.92~{}m$, and is located in the German Bight, near the mouth of the river Elbe close to Hamburg (Germany). This area is a hot spot for storm surges, and also includes the largest storm surge range in the hindcast of $5.65~{}m$, as seen in Figure 8c. Generally, the southeastern part of the North Sea is an area of large storm surge activity, as is also illustrated by the large values of standard deviation for the DE, NL and DK stations in Figure 3a.
@@
A 43-year long hindcast is not sufficiently long to provide robust statistics of extreme water levels by direct analysis and extraction of extremes from the model results. We have therefore performed a statistical analysis based on annual (block) maxima of the hindcast dataset and calculated the 100-year Return Value Estimates (RVE) of minimum and maximum storm surge for the model domain by fitting a Gumbel distribution (Gumbel, 1958) to the annual minima and maxima, respectively. As shown by Coles (2001), the cumulative distribution function (CDF) of the block maxima formed from a random sequence of independent variables and identically distributed (IID) variables will follow the Generalized Extreme Value (GEV) distribution,
@@ 4.3.1 The 1987 Oslofjord storm surge
As the low pressure system moved further east, and the wind direction continued backing further left, the storm surge signal moved further through the North Sea up along the western coast of Denmark and into the Skagerrak area as a Kelvin wave (see Figure 12). Even though not directly affected by the weather system itself, the Oslofjord experienced large fluctuations in water level due to the Kelvin wave, with variations in water level due to storm surge in the inner part of the Oslofjord in the range of up to ±$\pm 1~{}m$ in less than 24 hours. This can be seen in the lower panel in Figure 12, displaying the observed and modeled storm surge for the station Viker (NO_VIK) at the entrance to the Oslofjord. Both the amplitude of the surge, and the values of MAE and RMSE, is significantly reduced compared to the two panels above for Cuxhaven and Esbjerg (DK_6401). Correlation between hindcast and observations for the three stations for the time period shown is almost identical at $0.93$ and $0.94$.
@@ 5 Summary and concluding remarks
The NORA-Surge hindcast is an open and freely available storm surge hindcast [^4] Available from https://thredds.met.no/thredds/catalog/stormrisk/catalog.html. based on a barotropic setup of the ROMS ocean model with $4$ km horizontal resolution covering the areas of the North Sea, the Norwegian Sea and the Barents Sea. It includes the entire coast of the United Kingdom, the Netherlands and Norway (including Svalbard), in addition there is partial coverage of the coast of Denmark and western coast of Sweden. The temporal coverage of the hindcast is more than four decades, covering the time period from 1979 throughout 2022.
The hindcast has been extensively validated against an observational dataset of quality checked observations of water level from more than 90 coastal and offshore locations, as described in Section 2. Comparison of the hindcast with all the observational data yields average values of MAE and RMSE of $9.7$ and $12.4$ cm respectively, scaled values for MAE/$\sigma$ of $0.42$ and RMSE/$\sigma$ of $0.54$ standard deviations (see Table 1 and 2), and correlation coefficients for each station primarily in the range of $0.80-0.93$ (see Figure 6). The absolute error varies between the stations, and has larger values for some geographical areas than others. The areas with the largest amplitudes in storm surge are also associated with the largest errors (see Figure 3 and Table 2). However, as shown in the Taylor diagram in Figure 6, the relative errors for the stations (with a few exceptions) are quite similar. There is a slight clustering of the stations from different countries, which reflects geographical and bathymetry specificities of the coastline depending on the location, but all the clusters have at least some overlap with each others.
Statistics of extreme values of minimum, maximum and range for storm surge within the hindcast are calculated and presented in the maps in Figure 8. The minimum storm surge value within the hindcast is $-2.60$ m, the maximum is $3.92$ m and the largest range of storm surge is $5.65$ m. All of which are found in the southern part of the North Sea. Also, a return value estimation using the GEV distribution has been carried out based on the hindcast dataset, and the 100-year return value estimates of minima and maxima of water level for the entire hindcast model domain are presented in the maps in Figure 9. The minimum and maximum return values for storm surge are $-2.75$ m and $3.98$ m, respectively, also located in the southern North Sea.
@@
The hindcast generally validates well against a broad range of water level measurements in the North Sea. However, despite mostly capturing the range and also upper percentiles, a number of improvements can be made. As has been mentioned, increased horizontal resolution with more detailed bottom bathymetry and a more realistic coastline, would likely reduce some errors in the hindcast. In addition, the inclusion of tides should be considered to improve the quality in regions with large tidal amplitudes and shallow water depths where surge-tide interaction is expected to be important. In very shallow areas, wave setup can also add to the water level in storm conditions (Stockdon et al., 2006; Melet et al., 2018). Finally, we expect the atmospheric forcing in the storm surge model could be improved by including wave effects in the atmospheric momentum fluxes (Staneva et al., 2017; Bonaduce et al., 2020).
For users of the hindcast dataset, we encourage to subtract the hindcast mean storm surge water level for each grid point from the data before use, as we have done. Additionally, various methods for correction could be explored and applied to the hindcast dataset to further improve the quality. But this was not within the scope of the present work.
@@ Acknowledgments
The license agreements of the water level observations we gathered do not allow to redistribute our dataset of in-situ observations. However, the dataset can be shared upon reasonable request and collaboration by taking contact with the corresponding author of this study. Moreover, the list of data sources presented in Table 3 can be used to retrieve again the data. Though this involves a significant amount of work, we provide a series of code repositories that can be used as a basis or inspiration for building the aggregated dataset:
Appendix B: Observation data collection, quality control, and computation of the sea elevation residual
Prefix Source Data request URL
NO Kartverket - Norwegian Mapping Authority https://api.sehavniva.no/tideapi_en.html
WL Norwegian Meteorological Institute, oil platforms sensors MetNo internal data: address data request to ocean@met.no
SW Swedish Meteorological and Hydrological Institute https://www.smhi.se/data/oceanografi/ladda-ner-oceanografiska-observationer/#param=sealevelrh2000,stations=all
DK Danish Coastal Authority https://kyst.dk/hav-og-anlaeg/maalinger-og-data/vandstandsmaalinger
DE BFG - Federal Institute of Hydrology https://www.bafg.de/EN/06_Info_Service/01_WaterLevels/waterlevels_node.html
NL Rijkswaterstaat environment https://waterinfo.rws.nl/#/publiek/waterhoogte
UK BODC - British Oceanographic Data Center https://www.bodc.ac.uk/data/hosted_data_systems/sea_level/uk_tide_gauge_network/processed/
@@
Figure 14: Time series of observed total water level from station DK_4203 showing an example of different types of flagged values from the SaQC analysis.
@@
  • For each data source, a custom parser (or even several parsers, as some of the data sources have format differences between different time periods and stations) was developed to convert the different data formats used into a list of tuples: $(time\_of\_observation,water\_level\_observation,quality\_flag)$. From there, a time series of the raw collected observation data is established for each station (note that the quality_flag is available on a data source dependent basis).
  • The time series for each station is, at this stage, a series of observation tuples, which temporal resolution varies between stations (and for some stations, with time), from 30 minutes (coarsest observations assimilated in our dataset) to 10 minutes (finest observations available for some stations). We perform a simple quality checking, and we first reject data that have either a faulty quality flag (when provided), or that have clearly faulty values (in particular unrealistic absolute value, such as for example values of 999.9 reported by some sensors as an error code; moreover, some instrument malfunctions can also result in readings that clearly stand out of the typical data range but with a less drastic absolute value, and these are detected later on by our quality control procedure, see illustration in Figure 14). The amount of such data points rejected varies from station to station, but typically less than 5% of the data are flagged as such in the stations reported here. We then perform interpolation of these data onto a common time basis, with 10 minutes resolution (corresponding to UTC minutes equal to 0 modulo 10).
At this stage, consistent time series on a common time basis are obtained for all the stations. We then perform a second quality check, that is based on human investigation. In particular, the time series for each station was visually inspected for problems such as large average trends in the data (this happens, for example, for the oil platforms, that are standing on the seafloor and subject to seabed subsidence), and change of the reference levels following sensor replacement, maintenance, or re-positioning. These errors are detected by directly inspecting the signals, and the time series are then corrected into 0-mean data, either by linearly detrending the data between sensor re-positioning (at the oil station locations where subsidence takes place), or by computing the mean water elevation to subtract on segments of data in between sensor servicing when no subsidence is present.
@@
At each station location, the full quality controlled sea elevation signal is then used to perform a standard harmonic tidal analysis using the pytide software tool, as described in the main body of the manuscript. This allows to obtain a best estimate for the weather-independent astronomical tide contribution to the sea level. Following this, we use the residual obtained by subtracting the harmonic tide contribution from the sea elevation observation to estimate the weather sea level surge component. This is illustrated, for a typically representative station (here, the Mausund NO station, though results obtained at other stations look similar), in Fig. 15. As visible there, the harmonic tide analysis captures the tide contribution well, and clear low-frequency weather effects corresponding to storms passing by and the associated storm surges are visible in the residual.
Figure 15: Illustration of the quality controlled observations, the harmonic tide analysis computed from the full duration of the observations timeseries using pytide, and the residual signal obtained by subtracting the former two, during part of the 2020 winter at the Mausund NO station. The figure is typically representative of what is obtained at other locations. The harmonic tide analysis captures well the tidal component of the sea level. The residual signal contains clear signatures of storms passing by the station, and the associated storm surges.